### v1

# Resolving the Ambiguity of Complete-to-Partial Point Cloud Registration for Image-Guided Liver Surgery with Patches-to-Partial Matching

Abstract
In image-guided liver surgery, the initial rigid alignment between preoperative and intraoperative data, often represented as point clouds, is crucial for providing sub-surface information from preoperative CT/MRI images to the surgeon during the procedure. Currently, this alignment is typically performed using semi-automatic methods, which, while effective to some extent, are prone to errors that demand manual correction. Point cloud correspondence-based registration methods are promising to serve as a fully automatic solution. However, they may struggle in scenarios with limited intraoperative surface visibility, a common challenge in liver surgery, particularly in laparoscopic procedures, which we refer to as complete-to-partial ambiguity. We first illustrate this ambiguity by evaluating the performance of state-of-the-art learning-based point cloud registration methods on our carefully constructed in silico and in vitro datasets. Then, we propose a patches-to-partial matching strategy as a plug-and-play module to resolve the ambiguity, which can be seamlessly integrated into learning-based registration methods without disrupting their end-to-end structure. It has proven effective and efficient in improving registration performance for cases with limited intraoperative visibility. The constructed benchmark and the proposed module establish a solid foundation for advancing applications of point cloud correspondence-based registration methods in image-guided liver surgery. Code and datasets will be released.
Index Terms:
Image-guided liver surgery, point cloud matching, pre- to intraoperative rigid registration.
I Introduction
In liver surgery, preoperative imaging techniques such as Computed Tomography (CT) and Magnetic Resonance Imaging (MRI) provide detailed information about critical anatomical structures, such as blood vessels and tumors. However, this information is not usually available during surgery due to the large footprint of the equipment, its immobility, and its prohibitively high cost. Consequently, intraoperative imaging is most commonly limited to lower resolution or partial information covering only a portion of the visible organ from modalities including intraoperative ultrasound [1, 2], cone-beam CT (CBCT) [3, 4], and optical cameras [5, 6, 7, 8]. To inform the surgeon of subsurface information from preoperative CT/MRI during surgery, a registration process is involved to compute the transformation between the preoperative and intraoperative data in image-guided surgery (IGS) systems, which is generally divided into two steps.
Step 1 is an initial rigid registration to resolve the problem that preoperative and intraoperative data are acquired in different coordinate systems. However, current IGS systems often perform this alignment semi-automatically [9], bringing considerable time consumption and susceptibility to errors [10].
Step 2 involves non-rigid registration methods to estimate tissue deformations, which depend on an accurate initial rigid registration.
Figure 1: Comparison of widely used public point registration datasets in computer vision (3DMatch, 4DMatch, ModelNet40) and our liver registration dataset. In each example, the source point clouds (blue) are aligned with the target point clouds (red). 3DMatch, 4DMatch, and ModelNet40 focus on partial-to-partial cases. The liver registration dataset is distinct from these datasets, designed under the complete-to-partial scenarios.
Registration methods that adapt to available intraoperative information can be categorized into intensity-based, 3D-to-2D, and 3D-to-3D methods. This paper addresses the initial rigid registration problem in a 3D-to-3D context, where both the preoperative and intraoperative data are represented as 3D point clouds. The preoperative point cloud is often derived from CT or MRI scans, while the intraoperative point cloud is often captured using 3D sensors or reconstructed from endoscopic video images via 3D reconstruction. For clarity, we refer to the point cloud representing the entire liver from the preoperative surface as the source point cloud and the partial intraoperative point cloud as the target point cloud.
Learning-based point cloud registration methods, particularly correspondence-based approaches, have shown promising results [11, 12, 13, 14] to solve the limitations of semi-automatic methods. However, applying these methods, designed for natural datasets, to liver surgery is challenging due to two reasons:
[…]
As a further challenge, the smooth surface of the liver lacks strong visual key points that could assist with matching partial point clouds and complicates current approaches to accurately align the visible intraoperative portion with the complete preoperative liver model (i.e., see Fig. 1 (d), where the target point cloud has a limited visible surface area). Besides the confusion associated with matching to the top or bottom of the liver, there is also ambiguity when registering a small point cloud to regions to regions adjacent to the liver dome.
As such, we argue that current learning-based point cloud registration methods are insufficient to handle the complete-to-partial ambiguity that frequently arises in LLS.
II. Substantial data requirements. The success of learning-based point cloud registration methods is heavily based on scaled-up datasets. However, this is particularly challenging for liver surgery as collecting extensive in vivo or in vitro datasets is expensive, complex, and time-consuming. A common practice to capture the preoperative and intraoperative configurations of the liver is to embed fiducial markers. However, this process may cause damage to tissues or phantoms. An alternative is to use simulated data for training and then apply the model to real-world in vivo or in vitro data [18]. While promising, there is currently a scarcity of large, publicly accessible simulated datasets, and the availability of real-world testing datasets remains severely constraining.
I-A Contribution
To address the first challenge, we propose a novel, learnable-parameter-free approach that integrates seamlessly with correspondence-based point cloud registration methods to mitigate complete-to-partial ambiguity (see §III-B). Our approach significantly improves low-visibility performance with negligible computational overhead. The core concept is to transform the complete-to-partial point cloud registration problem into a subdivided matching problem between the partial point cloud and multiple patches extracted from the complete cloud, which we refer to as patches-to-partial (P2P) matching.
In the P2P, we first select the most likely visible points from the source point cloud. Then, we generate several candidate patches of approximately the same size as the target point cloud. Finally, we match the target point cloud to these candidate patches and aggregate a robust, rigid point cloud transformation by selecting the best estimate across candidates. This module only involves feature resampling and rematching, allowing it to be parallelized for fast execution. Additionally, it is fully differentiable, making it suitable for integration into different correspondence-based point cloud and learning-based registration pipelines.
To address the second challenge, we developed a large simulated dataset comprising over $1,000$ liver models and $10,000$ simulations alongside phantom datasets (see §IV-A), providing valuable resources for benchmarking and future advancing research in the liver registration area.
We experimentally demonstrate that the current state-of-the-art correspondence-based registration methods are not sufficient when dealing with complete-to-partial ambiguity. We then show the effectiveness of our proposed module in dealing with low-visibility cases after seamless plug-and-play integration into two promising methods. Without any adjustment to loss functions, additional learnable modules, outlier rejection methods, etc., our approach significantly improves complete-to-partial registration.
II Related work
II-A Traditional point cloud registration
Iterative Closest Point (ICP) [19] and its variants are commonly used in semi-automatic registration processes. ICP iteratively finds the closest points between two point clouds and estimates the rigid transformation using singular value decomposition (SVD) [20]. Though intuitive, ICP is sensitive to initial alignment. Globally optimal ICP (GO-ICP) [21] poses an alternative approach by introducing a global search strategy to improve robustness. However, it then becomes computationally expensive and does not always guarantee an optimal solution.
II-B Learning-based point cloud registration
Learning-based point cloud registration methods can be broadly categorized into correspondence-based and direct registration approaches. Correspondence-based methods [12, 13] first establish corresponding point pairs between two point clouds, then compute the optimal transformation based on these correspondences. Direct registration methods [22, 23], on the other hand, directly regress the rigid transformation from global features. While these methods show promising results on synthetic, object-centric point clouds [15], they usually struggle with low overlap ratios, limited transformation ranges, and larger-scale scenes [12, 24]. Currently, the state-of-the-art techniques assessed on publicly available benchmarking datasets [11, 12, 13] are correspondence-based.
Representative correspondence-based methods, such as 3DMatch [11], Predator [12], and Lepard [13], introduce techniques like Siamese networks, attention modules, and positional encoding to improve point cloud matching. These techniques and their variants have since become standard practice in correspondence-based methods. While most of these methods focus directly on fine-scale correspondences, some, such as those reported in [14, 25, 26], adopt a point-to-node grouping strategy, where coarse-scale matches guide fine-scale correspondences within patches. However, whether this strategy performs well on datasets with complete-to-partial ambiguity remains unclear. Thus, a representative method [26] using the strategy is included in the experiments.
Early correspondence-based methods [11, 12] rely on RANSAC, a widely known robust estimator that iteratively selects subsets of correspondences to find the best transformation. However, RANSAC often suffers from slow convergence [27] and is unsuitable for end-to-end learning schemes. Consequently, recent methods aim to achieve comparable results without relying on RANSAC. For instance, RegTr [24] directly predicts keypoint displacements, while RoITr [26] uses a local-to-global registration approach to eliminate the need for RANSAC.
Overall, correspondence-based methods have a strong foundation in several public benchmarks for natural datasets. Still, they are not designed to handle the complete-to-partial cases encountered in liver surgery.
II-C 3D-3D Initial Rigid Registration in Liver Surgery
Automatic correspondence methods using hand-crafted feature descriptors have been explored by Dos Santos et al. [28] and Robu et al. [29]. However, these methods have limited accuracy and often require additional pruning.
Several learning-based methods [30, 31, 32] have been proposed for 3D-3D liver registration, yielding promising results. However, these outcomes should be interpreted with caution. For example, the simulation dataset used in [31] assumes one-to-one correspondence, as both the source and target point clouds are derived from the same mesh without post-processing. Consequently, models pre-trained or fine-tuned on these datasets may struggle when applied to real-world data. Furthermore, Guan et al. [30] and Zhang et al. [32] treat the intraoperative CT surface as the preoperative surface in the porcine dataset [33], resulting in source and target point clouds with no deformation.
In contrast, we carefully post-process our simulations to include soft tissue deformations, eliminate one-to-one correspondences, and remove rigid components from the simulated deformations. We evaluate the selected registration methods on both our simulation and phantom datasets, ensuring a more rigorous and realistic assessment.
III Methodology
Figure 2: Illustration of the general paradigm of learning-based correspondence registration methods (top) and our plug-and-play patches-to-partial (P2P) module (bottom). Our proposed module generates candidate patches, samples their point-wise features from a learning-based correspondence registration method, and performs feature matching and rigid transformation estimation for each patch with the target point cloud. Finally, a candidate selection rule determines the optimal rigid transformation. We present the case where the patch number $k$ is set to 5.
This section introduces our proposed patches-to-partial strategy as a plug-and-play module that can be seamlessly integrated into current learning-based correspondence registration methods. We begin by presenting the general paradigm of learning-based correspondence registration methods (see §III-A), followed by a detailed module explanation (see §III-B). The overall framework is illustrated in Fig. 2.
III-A Preliminaries
Given a source point cloud $\mathbf{S}\in\mathbb{R}^{N\times 3}$ (preoperative) and a target point cloud $\mathbf{T}\in\mathbb{R}^{M\times 3}$ (intraoperative) as inputs, correspondence-based registration methods output their respective point-wise features, $\mathbf{x}^{\mathbf{S}}\in\mathbb{R}^{N\times d}$ and $\mathbf{x}^{\mathbf{T}}\in\mathbb{R}^{M\times d}$, along with the rigid transformation from $\mathbf{S}$ to $\mathbf{T}$. Here, $N$ and $M$ represent the number of participating points in the source and target point clouds, respectively, and $d$ is the feature dimension. These registration methods generally consist of four key components: Feature extraction network, Matching module, Rigid transformation estimation module, and Loss functions.
Feature extraction networks extract point-wise feature descriptors from both the source and target point clouds. These networks typically use a point cloud feature learning backbone, such as PointNet++ [34], KPConv [35], or Point Transformer [36], which usually involves a U-Net-like structure that gradually down-samples the input point clouds to extract features, followed by upsampling to restore the resolution of the original points while preserving the extracted features.
Attention mechanisms, which have emerged as a pivotal component for feature understanding in various domains [12, 13, 26], are frequently incorporated within the feature learning backbone, particularly in the bottleneck, to empower the interaction between local geometric features towards global geometry-aware ones. This interaction can occur either within the same point cloud (self-attention) or across different point clouds (cross-attention), depending on the layer used.
Matching modules then use either Sinkhorn optimal transport algorithm [37, 26] or dual-softmax operation [38, 13] to match the source and target feature descriptors, which output correspondences between source and target point clouds and their confidences.
Rigid transformation estimation modules further utilize the estimated correspondences and confidences from the matching module and calculate the rigid transformation, using the SVD or weighted SVD [39, 19]. If the estimated correspondences contain a large portion of outliers, the weighted SVD tends to predict an inaccurate rigid transformation. In this case, additional outlier rejection methods are involved in sampling inlier correspondence (e.g., RANSAC).
Loss functions consist of one or more types, such as classification-based loss[13] (e.g., cross-entropy with dual-softmax or optimal transport), feature metric-based loss[12, 26] (e.g., circle loss), and displacement-based loss for predicted matches or estimated rigid transformation accuracy in 3D space.
III-B Transferring complete-to-partial into patches-to-partial matching
The smoothly shaped surface of the liver produces a challenge for point cloud matching wherein the partial target point cloud may incorrectly match well with multiple candidate regions within the source point cloud due to shallow variations in surface curvature. Although correspondence-based registration methods are well-established in several public natural benchmarks, they are not designed to address the complete-to-partial ambiguity inherent in this scenario.
This challenge is similar to a Jigsaw puzzle, where a player tries different candidate spaces to find the correct fit. Inspired by this analogy, we propose converting the complete-to-partial point cloud matching problem into a patches-to-partial (P2P) matching process. By generating several candidate patches from the source point cloud, each roughly the size of the target, and matching them to the target, we can naturally solve the complete-to-partial ambiguity. The questions thus become clear:  How do we efficiently generate candidate patches from the source point cloud?  How do we select the best transformation?
For consistency in the presentation, we use the point-wise features $\mathbf{x}^{\mathbf{S}}$ and $\mathbf{x}^{\mathbf{T}}$, along with the estimated transformation ${\mathbf{R}_{c},\mathbf{t}_{c}}$ from existing correspondence-based registration methods, as our starting point to address these questions and validate this approach.
III-B1 Visible Source Point Estimation
In response to , a straightforward approach is to recognize each point from the source point cloud as a patch center and select its nearest neighbors, matching the number of points in the target cloud to generate candidate patches. However, this approach requires impractical computational memory. A more efficient approach would be to first narrow down the candidate points before generating these patches, which this part seeks to present rigorously.
To narrow down the candidate points, we first design a method to select the top candidates from the source point cloud based on a visibility score, which we define according to a score matrix following [40, 13] as $S=\mathbf{x}^{\mathbf{S}}(\mathbf{x}^{\mathbf{T}})^{\top}\in\mathbb{R}^{N\times M}$, where $S$ represents the cosine similarity between source and target features. We then convert this score matrix into visibility scores for the source point cloud by summing over the target dimension: $S_{vis}=\sum_{j=1}^{M}S_{i,j}$. Finally, we select the top $M$ source points with the highest visibility scores to form the visible source point cloud.
III-B2 Patch Candidate Generation
The visible source point cloud contains both correct and uncertain points due to complete-to-partial ambiguity. The points are distributed across a region, and it is assumed that matching a patch within this region to the target point cloud facilitates a more precise estimation of rigid transformation compared to utilizing the entire point cloud. Consequently, our next step is to generate a few patches from the region. To do this, we apply farthest point sampling (FPS) [16] to the visible source point cloud, selecting $k$ points to serve as patch nodes.
III-B3 Patch Sampling and Feature Computation
The 3D locations and features of candidate patches are then sampled and generated from the nearest neighbors of these patch nodes, with the same number of points as the target point cloud. We denote the features and 3D locations of the $k$-th source patch as $\mathbf{x}^{\mathbf{S}_{k}}\in\mathbb{R}^{M\times d}$ and $\mathbf{S}_{k}\in\mathbb{R}^{M\times 3}$, respectively.
III-B4 Patch-to-Target Matching
Each patch candidate generated in the previous step is then matched to the target point cloud to establish correspondences via rigid transformation estimation. Taking one patch as an example, we first compute the score matrix $S_{k}=\mathbf{x}^{\mathbf{S}_{k}}(\mathbf{x}^{\mathbf{T}})^{\top}\in\mathbb{R}^{N\times M}$, and apply the dual-softmax operation [40, 13, 38]:
$M_{k}(i,j)=\text{Softmax}(\mathcal{S}_{k}(i,:))\cdot\text{Softmax}(\mathcal{S}_{k}(:,j)),$ (1)
where the dual-softmax operation converts $S_{k}$ into a confidence matrix $M_{k}$. Matches are selected from $M_{k}$ using the mutual nearest neighbor criterion, where a pair of indices $(i,j)$ is chosen if its confidence value $M_{k}(i,j)$ is the maximum in both $\mathcal{S}_{k}(i,\cdot)$ and $\mathcal{S}_{k}(\cdot,j)$.
Once the correspondence set $\mathcal{C}_{k}$ is obtained, the weighted SVD is applied to compute the rigid transformations $\mathbf{R}_{k}$ and $\mathbf{t}_{k}$. The transformations estimated from each patch are inputs to the next step.
III-B5 Optimal Candidate Selection
A candidate selection procedure is needed in our design to select the most likely rigid transformation from the patches, answering . One possible solution for candidate selection is to apply the inlier number selection rule used in the point-to-node strategy [25, 26]. This approach selects the best transformation by choosing the one that results in the most inlier matches:
$\mathbf{R},\mathbf{t}=\max_{\mathbf{R}_{i},\mathbf{t}_{i}}\sum\nolimits_{(p_{j}^{S},p_{j}^{T})\in\mathcal{C}_{all}}\left[\lVert\mathbf{R}_{j}\cdot p_{j}^{S}+\mathbf{t}_{j}-p_{j}^{T}\rVert_{2}^{2}<\tau\right],$ (2)
where the Iverson bracket []delimited-[][$\cdot$][ ⋅ ] evaluates whether the condition holds, $\tau$ is the acceptance radius. $\mathcal{C}_{all}$ combines all correspondences from the original output and all $\mathcal{C}_{k}$ from the previous step. $p_{j}^{S}$ and $p_{j}^{T}$ are 3D points within the correspondences.
The transformations $\mathbf{R}_{i}$ and $\mathbf{t}_{i}$ are from the original output and all candidate transformations ${\mathbf{R}_{k},\mathbf{t}_{k}}$. However, this inlier-based selection favors more feature-matching evaluation metrics (e.g., inlier ratio [12]), while our goal is to select the best rigid transformation. In light of this view, we propose a straightforward selection rule by directly calculating the mean closest point distance between the transformed source point cloud, $\mathbf{R}_{i}\cdot x^{S}+\mathbf{t}_{i}$, and the target point cloud $x^{T}$ as:
$\mathbf{R},\mathbf{t}=\min_{\mathbf{R}_{i},\mathbf{t}_{i}}\sum_{j=1}^{M}\mathbf{Dist}_{min}(\mathbf{R}_{i}\cdot x^{S}+\mathbf{t}_{i},x_{j}^{T}),$ (3)
where $\mathbf{Dist}_{min}$ is the closest point operator that computes the distance from $x_{j}^{T}$ to the transformed source point $\mathbf{R}_{i}\cdot x^{S}+\mathbf{t}_{i}$.
Figure 3: Illustration of the in silico phantom generation process. Source and target point clouds/ meshes are shown in blue and red, respectively.
IV Experiments
IV-A Datasets and Evaluation
We use the in silico dataset for both training and testing, while the in vitro dataset is used exclusively for testing. The same procedure (see §IV-B1) is applied to preprocess the input data for the baseline and the proposed methods (see §IV-B2).
IV-A1 In silico phantom dataset
Fig. 3 illustrates the generation process of the in silico dataset. We utilized CT scan volumes from the training dataset of the Medical Segmentation Decathlon [41] to generate liver models represented by $131$ triangular meshes. These meshes were extracted using the marching cubes algorithm from the scikit-image Python library [42]. To enhance the quality of the mesh, we applied a Laplacian smoothing kernel and used vertex clustering with a voxel size of $2$ mm from the Open3D Python library [43]. Finally, we used Tetgen [44] to generate volumetric mesh from the surface meshes. The liver models were split into $120$ models for training and 11 models for testing. Each liver model was randomly scaled in the x, y, and z dimensions to augment the data. The isotropic scale factors were uniformly generated from the range [$0.5$, $1$]. This process generated $10$ additional liver models for each original liver, resulting in $1,320$ liver models for training and $121$ liver models for testing.
For each liver model, we generated $10$ deformations using the pipeline proposed in [18] via a finite element model and used its default parameters. Some simulated deformations feature rigid transformations instead of deformation when the randomly generated zero boundary condition is small. To eliminate rigid transformation, we further align the undeformed and deformed models with the SVD using the volumetric vertices as fiducial markers. After the alignment, the average root mean square error of the fiducial markers is $3.37$ mm, with a maximum value of $11.38$ mm, which is similar to the deformation range in Sparse Non-rigid Registration Challenge dataset [45].
For both training and testing datasets, we followed the previous works [12, 13, 31] to perform random cropping and rigid transformations on the deformed surface for each deformation. We calculate the visibility ratio as $M/N$. The visibility of the target point clouds was between $0.2$ and $1.0$. Regarding the rigid transformations, we sample rotations according to three Euler angle rotations in the range $[0,2\pi]$ and translations in the range [$-100$, $100$] mm.
For the testing dataset, we generated five target point clouds for each deformation. Consequently, the testing dataset contained $121$ source point clouds with $6,050$ target point clouds.
The training dataset comprised $1,320$ liver models with $13,200$ deformed liver models after the deformation simulation. A specialized training strategy will be introduced later in IV-B to utilize the training dataset.
As illustrated in Fig. 3, the volumetric vertices are assumed to be fiducial markers for measuring registration accuracy.
IV-A2 In vitro phantoms
We used the phantom dataset constructed in [46] as a baseline to create another testing dataset to prove the generality of our proposed method. The phantom dataset contains four undeformed (source) and deformed (target) models with known embedded fiducial markers as validation points to compute the target registration errors. CT scans of both undeformed and deformed phantoms were acquired, and the ground truth surface mesh and fiducial landmarks were manually segmented from the CT scans. We generated $200$ target point clouds from the complete target model surface, producing $800$ testing samples.
IV-A3 Evaluation
A large number of evaluation metrics [12, 13, 23] exist to measure point cloud registration. However, these metrics rely on ground truth rigid transformations or one-to-one correspondence, which do not apply to our settings. Also, they do not provide insight into subsurface registration errors.
Here, we use the root mean square target registration error (RMS-TRE) to measure the final outcome of the registration methods on the above datasets. RMS-TRE offers an unbiased assessment of registration accuracy across the liver, including fiducial markers underneath the surface, and directly reflects the precision required when overlaying preoperative data onto intraoperative data. It takes the form as:
[…]
where $\mathbf{Y}$ and $\mathbf{X}$ represent the locations of fiducial markers in the source and target, respectively, and $N_{f}$ is the number of fiducial markers.
IV-B Implementations
IV-B1 Preprocessing
The preprocessing step is crucial as the point cloud feature learning models typically require a consistent point cloud density to extract features effectively. We begin by centering the source point cloud by subtracting its centroid, then scale it to fit within a unit sphere using the maximum distance from the origin. The target point cloud is processed similarly but scaled using the same factor as the source. Both point clouds are then voxelized with a voxel size of $0.04$.
IV-B2 Baselines
We used the official implementations of the following methods: (1). Go-ICP [21], a traditional method that globally searches for a minimizing condition across SE(3). (2). RegTr [24], an end-to-end point cloud correspondence network that predicts a final set of correspondences directly rather than matching source and target features. (3). RoITr [26], a state-of-the-art method that utilizes a point-to-node strategy. (4). Lepard [13], another state-of-the-art method, excelling in rigid and non-rigid point cloud registration datasets. (5). LiverMatch [31], an available correspondence-based point cloud registration method specifically developed for liver surgery. RegTr [24], Lepard [13], and LiverMatch [31] leverage the power of KPConv[35] in their point cloud feature learning. The initial cell size in KPConv, which controls the downsampling rate, is set to $0.02$ to enable gradual downsampling on our datasets. All the learning-based methods are implemented in PyTorch [47].
IV-B3 Proposed module
The only hyperparameter of the proposed module is the number of candidate patches, K, whose default value is set to $5$. In §V-C, a sensitivity study is conducted to evaluate the impact of this hyperparameter on the downstream performance of the registration methods.
IV-B4 Procrustes
In addition to the baseline methods, we also include Procrustes, which uses SVD with ground truth fiducial markers to estimate the rigid registration. This method serves as a reference for deformation and potentially offers the best rigid registration that baseline methods can achieve.
IV-B5 Statistical hypothesis test
The Wilcoxon Rank Sum Test at a significance level of $\alpha=0.05$ is used to assess whether integrating the proposed module with a baseline method significantly improves registration results compared to the original baseline performance.
IV-B6 Training
We used a special training strategy that fully exploits the in silico training dataset by dynamically selecting source and target point cloud pairs during training. Specifically, two source and target point clouds are randomly chosen from a liver and its deformations within the training dataset. Random cropping and rigid transformations are then applied to the target point cloud, as described in §IV-A1. Some baselines required either rigid transformation or correspondences between source and target point cloud. The required rigid transformation is obtained from Procrustes, and the correspondences are determined using nearest neighbor searching between the source and target point clouds within a radius of $0.04$ after applying the Procrustes. For all methods, the maximum number of training epochs is set to $150$, and the batch size is set to $1$. Training is performed on an NVIDIA A100 GPU, with Lepard taking approximately three days to train, while other methods require about one day.
V Results
V-A In Silico Phantom Validation
TABLE I: Comparison of registration errors on the In Silico phantom dataset across different visibility ratios. The average RMS-TRE ± standard deviation is reported in millimeters.
Visibility ratio $[0.2,0.3)$ $[0.3,0.4)$ $[0.4,0.5)$ $[0.5,0.6)$ $[0.6,0.7)$ $[0.7,0.8)$ $[0.8,0.9)$ $[0.9,1]$
Procrustes 3.19 $\pm$ 1.93 3.44 $\pm$ 2.16 3.41 $\pm$ 2.13 3.42 $\pm$ 2.19 3.38 $\pm$ 2.17 3.41 $\pm$ 2.17 3.39 $\pm$ 2.01 3.31 $\pm$ 2.08
GO-ICP [21] 66.96 $\pm$ 17.2 66.98 $\pm$ 19.52 56.4 $\pm$ 23.52 55.61 $\pm$ 25.41 44.83 $\pm$ 29.27 33.56 $\pm$ 29.94 19.9 $\pm$ 24.21 5.42 $\pm$ 2.65
RoITr [26] 20.21 $\pm$ 25.12 10.77 $\pm$ 12.71 7.24 $\pm$ 8.13 5.6 $\pm$ 6.05 4.61 $\pm$ 3.57 4.19 $\pm$ 2.61 3.91 $\pm$ 2.36 3.75 $\pm$ 2.46
RegTr [24] 84.27 $\pm$ 29.46 69.89 $\pm$ 30.16 59.66 $\pm$ 30.45 48.08 $\pm$ 30.11 33.53 $\pm$ 24.36 20.24 $\pm$ 11.89 12.05 $\pm$ 4.16 5.82 $\pm$ 2.41
LiverMatch [31] 12.85 $\pm$ 13.87 8.09 $\pm$ 6.88 6.65 $\pm$ 5.15 5.65 $\pm$ 3.98 4.79 $\pm$ 2.82 4.24 $\pm$ 2.45 3.84 $\pm$ 2.09 3.56 $\pm$ 2.07
LiverMatch[31] + proposed P2P 8.74 $\pm$ 10.23 6.55 $\pm$ 5.54 5.5 $\pm$ 3.69 5.00 $\pm$ 3.30 4.40 $\pm$ 2.71 4.08 $\pm$ 2.46 3.78 $\pm$ 2.15 3.58 $\pm$ 2.16
Lepard [13] 9.47 $\pm$ 14.74 6.17 $\pm$ 4.83 5.22 $\pm$ 3.7 4.76 $\pm$ 4.33 4.20 $\pm$ 2.58 3.96 $\pm$ 2.43 3.74 $\pm$ 2.17 3.57 $\pm$ 2.22
Lepard [13] + proposed P2P 6.73 $\pm$ 5.96 5.63 $\pm$ 3.72 4.96 $\pm$ 3.28 4.60 $\pm$ 2.93 4.22 $\pm$ 2.68 4.00 $\pm$ 2.50 3.78 $\pm$ 2.22 3.59 $\pm$ 2.25
TABLE II: Registration errors on the in silico phantom dataset at different noise levels, within the visibility ratio range [0.2, 0.3). The average RMS-TRE ± standard deviation is reported in millimeters.
Noise level 0 mm 2 mm 4 mm
Procrustes 3.37 $\pm$ 2.11 3.37 $\pm$ 2.11 3.37 $\pm$ 2.11
LiverMatch [31] 12.85 $\pm$ 13.87 13.72 $\pm$ 15.48 21.85 $\pm$ 23.2
LiverMatch [31] + proposed P2P 8.74 $\pm$ 10.23 9.19 $\pm$ 9.48 14.58 $\pm$ 18.11
Lepard [13] 9.47 $\pm$ 14.74 10.07 $\pm$ 15.28 14.7 $\pm$ 21.02
Lepard [13] + proposed P2P 6.73 $\pm$ 5.96 7.22 $\pm$ 8.45 9.85 $\pm$ 14.88
Figure 4: Qualitative comparison of registration results on the in silico phantom dataset. The source and target point clouds are shown in blue and red, respectively.
V-A1 Performance against Baselines
We begin by evaluating the performance of the baseline methods under varying visibility ratios, with the quantitative results summarized in Table I.
At higher visibility ratios (i.e., visibility $\in[0.9,1.0]$), the baseline methods perform similarly to Procrustes, where rigid registration achieves the lowest errors, even in the presence of deformation. However, as the visibility ratio decreases, the performance of GO-ICP declines sharply, followed by RegTr, RoITr, LiverMatch, and Lepard. Among these, Lepard shows better robustness to changes in visibility, although its RMS-TRE still increases noticeably from 3.56 mm at visibility ratios $\in[0.9,1.0]$ to 9.47 mm at visibility ratios $\in[0.2,0.3]$
We then examine the results obtained by Lepard and LiverMatch with our proposed module, alongside those from RoITr quantitatively and qualitatively in Table I and Fig. 5, respectively. At the lowest visibility ratio range (i.e., visibility $\in[0.2,0.3]$), the registration errors for Lepard decrease from $9.47$ mm to $6.73$ mm, and for LiverMatch, from $12.85$ mm to $8.74$ mm following the integration of our proposed module. These improvements are statistically significant.
V-A2 Sensitivity to Noise
Given the promising results achieved by Lepard and LiverMatch with our proposed method, we further investigate their performance in the presence of noise and across groups with varying deformations in Table II. In our setting, noise is generated by sampling uniform random values from the range $[-0.5,0.5]$, scaling them by the noise level, and applying them to the target point cloud. When the noise level increases from $2$ to $4$ mm, the performances of LiverMatch and Lepard decline by $59.25\%$ and $43.92\%$, respectively. Our proposed method, on the other hand, helps both methods maintain lower RMS-TRE at these noise levels.
V-A3 Visual Assessment
We further visualize the results in Fig. 4. In the first row, we show a case with a high visibility ratio as baseline, where all methods produce reasonable results. However, in the second and last rows, when the visibility ratio is low, all methods struggle to align the source to the target point cloud properly. After incorporating our proposed module, both LiverMatch and Lepard are able to achieve reasonable results.
Figure 5: Comparison of registration errors from the in silico phantom dataset, with the average RMS-TRE values plotted against different visibility ratios using the bin centers from Table I.
V-B In Vitro Phantom Validation
TABLE III: Comparison of registration errors on the In Vitro phantom dataset across different visibility ratios. The average RMS-TRE ± standard deviation is reported in millimeters.
Visibility ratio $[0.2,0.3)$ $[0.3,0.4)$ $[0.4,0.5)$ $[0.5,0.6)$ $[0.6,0.7)$ $[0.7,0.8)$ $[0.8,0.9)$ [0.9,1)][0.9,1)][ 0.9 , 1 ) ]
Procrustes 3.66 $\pm$ 1.69 3.67 $\pm$ 1.78 3.91 $\pm$ 1.74 3.58 $\pm$ 1.8 3.43 $\pm$ 1.68 3.62 $\pm$ 1.7 3.77 $\pm$ 1.6 3.65 $\pm$ 1.91
GO-ICP [21] 87.15 $\pm$ 25.91 82.58 $\pm$ 25.45 75.34 $\pm$ 25.53 72.2 $\pm$ 30.03 61.89 $\pm$ 30.91 47.04 $\pm$ 36.66 21.16 $\pm$ 26.16 9.31 $\pm$ 19.08
RoITr [26] 33.8 $\pm$ 24.06 17.03 $\pm$ 17.2 12.67 $\pm$ 14.43 8.34 $\pm$ 9.33 5.7 $\pm$ 3.69 4.66 $\pm$ 2.21 4.39 $\pm$ 1.8 4.1 $\pm$ 2.09
RegTr [24] 66.25 $\pm$ 16.15 55.91 $\pm$ 19.19 45.02 $\pm$ 20.19 37.39 $\pm$ 21.67 29.74 $\pm$ 20.48 16.15 $\pm$ 13.07 8.76 $\pm$ 2.66 5.16 $\pm$ 1.8
LiverMatch [31] 20.54 $\pm$ 22.33 12.46 $\pm$ 12.87 8.26 $\pm$ 5.34 6.69 $\pm$ 4.29 5.26 $\pm$ 3.16 4.66 $\pm$ 2.58 4.18 $\pm$ 1.7 3.85 $\pm$ 1.85
LiverMatch [31] + proposed P2P 14.97 $\pm$ 18.73 9.21 $\pm$ 7.95 6.98 $\pm$ 4.51 5.66 $\pm$ 2.80 4.74 $\pm$ 3.03 4.5 $\pm$ 2.26 4.17 $\pm$ 1.74 3.83 $\pm$ 1.92
Lepard [13] 20.71 $\pm$ 23.39 10.32 $\pm$ 12.03 6.58 $\pm$ 4.33 5.3 $\pm$ 2.62 4.65 $\pm$ 3.09 4.24 $\pm$ 1.96 4.14 $\pm$ 1.72 3.86 $\pm$ 1.96
Lepard [13] + proposed P2P 12.45 $\pm$ 15.72 8.43 $\pm$ 9.64 6.07 $\pm$ 3.12 5.08 $\pm$ 2.58 4.45 $\pm$ 2.38 4.26 $\pm$ 2.01 4.19 $\pm$ 1.78 3.91 $\pm$ 2.02
Figure 6: Qualitative comparison of registration results on the in vitro phantom dataset. The source point cloud and its fiducial markers are shown in blue, while the target point cloud and its fiducial markers are shown in red, with fiducial markers displayed in larger sizes for visibility.
V-B1 Performance against Baselines
We further compare the baseline methods with and without our proposed module using the phantom datasets. The results are reported in Table III. As seen, the In Vitro results align with the findings from the In Silico experiments. GO-ICP, RegTr, and RoITr are more sensitive to visibility variations compared to LiverMatch and Lepard. After applying our module, the RMS-TRE values for LiverMatch and Lepard are reduced from $20.54$ mm to $14.97$ mm and from $20.71$ mm to $12.45$ mm, respectively. These improvements are statistically significant.
V-B2 Visual Assessment
For both the in silico and in vitro datasets, the target point clouds are automatically cropped. Additionally, we present qualitative results using manually cropped surfaces from [46], as shown in Fig. 6. In these manually cropped cases, the proposed method still significantly improves the registration results compared to Lepard. Overall, we demonstrate that the versatility of our proposed module enables seamless plug-and-play integration to improve the robustness and reliability of deep learning methods for rigid point cloud registration.
V-C Sensitivity and Ablation study
Figure 7: Sensitivity study on cluster number K, performed on the in vitro phantom dataset at different visibility ratio, using Lepard as the baseline for the proposed module.
TABLE IV: Ablation study on the selection method and candidate proposal, performed on the in vitro phantom dataset at a visibility range of [0.2, 0.3), using Lepard as the baseline for the proposed module. The average RMS-TRE ± standard deviation is reported in millimeters.
Selection rule Inlier number 14.14 $\pm$ 16.70
Closest point distance (default) 12.45 $\pm$ 15.72
Candidate proposal w/o 14.91 $\pm$ 19.06
w/ (default) 12.45 $\pm$ 15.72
Figure 8: Comparison of the proposed module with RANSAC ICP on the in vitro phantom dataset.
We further conduct comprehensive sensitivity and ablation studies on the phantom dataset, incorporating the proposed module with Lepard to evaluate its effectiveness.
Candidate patch number. The only hyperparameter in the proposed module is the candidate patch number, $K$. We vary this hyperparameter and present the registration results in Fig. 7, which shows a clear trend of decreasing registration error as the number of candidate patches increases. Statistical significance was observed between $K=4$ and $K=5$, but not between $K=5$ and $K=6$, suggesting that the default value of $K=5$ provides an optimal balance between computational complexity and accuracy.
Selection role. We further compare our closest distance-based selection introduced in Eq. 3, with the inlier number-based selection in Eq. 2, where the parameter $\tau$ is set to 0.05. As shown in Table IV, our approach proves to be more effective, resulting in fewer errors below 2 mm.
Patch candidate proposal. We also investigate the contribution of the candidate proposal to our proposed module. As shown in Table IV, omitting the patch candidate proposal results in a significant decline in performance, with a drop of approximately 2.5 mm in registration accuracy.
Comparison with RANSAC. Following previous studies [24, 26], we further compare the proposed module with RANSAC, a typical robust estimator used in point cloud registration. As discussed in Section II-B, the robust estimators do not fit the end-to-end learning scheme and are not designed to handle the complete-to-partial ambiguity.
Using the correspondence-based RANSAC ICP implementation from Open3D [43], with the hyperparameter for maximum correspondence distance set to 0.05, we found that our proposed module consistently outperforms RANSAC, particularly in low-visibility scenarios. RANSAC slightly worsened registration results in the visibility range $[0.4,1)$, likely due to its random sampling nature. The average running time of the proposed module is 0.04 seconds, compared to 1.06 seconds for RANSAC.
VI Discussion
While learning-based point cloud registration methods are grounded in a solid foundation in natural benchmarks, they face a unique challenge from the complete-to-partial ambiguity when applied to liver surgery scenarios, where only a partial surface is visible during an intraoperative operation.
To alleviate this ambiguity, we propose a patches-to-partial feature matching module to bridge the gap between methods and their applications in liver surgery. Also, as a comprehensive benchmark is missing in this field, we construct a benchmark consisting of in silico and in vitro datasets.
We then evaluate representative correspondence-based methods on the constructed datasets. While these methods achieve comparable registration errors in high-visibility cases, their performance deteriorates as the visibility ratio decreases. Among them, RegTr is particularly sensitive to reduced visibility, likely due to its unique design. RegTr directly predicts the displacements of source and target key points as correspondences, rather than matching features between the two point clouds. This approach works well when the source and target point clouds are of similar scale, but in cases of complete-to-partial ambiguity, predicting key point displacements is less robust due to the lack of global context from the target point cloud. RoITr, a method based on the node-to-group strategy, also shows sensitivity to low-visibility cases. Its node-to-group strategy is not designed to handle complete-to-partial ambiguity but rather aims to find more correspondences through a coarse-to-fine search. In comparison with RegTr and RoITr, LiverMatch and Lepard are less sensitive to decreasing visibility ratios. Both methods employ a U-Net-like structure and match features between the source and target point clouds at the final output, indicating that this simple pipeline may be more versatile. Despite showing promising results from LiverMatch and Lepard, their performances are still not satisfying in low-visibility cases. In such instances, after applying our module, we observe significant improvements $w.r.t.$ the registration performance (see Table I, III), having merits in:
i) Our proposed module alleviates the complete-to-partial ambiguity by converting it to patches-to-partial matching. Our module only has one hyperparameter, $K$, whose physical meaning controls the number of proposed patches. From the parameter-sensitivity study shown in Fig. 7, increasing $K$ will improve the registration by expanding a fully parallelizable search space, and the default value $K=5$ yields satisfactory results. In addition, we conducted an ablation study on our candidate proposal, and the closest point distance-based selection rule effectively contributes to the whole strategy.
ii) Our proposed module can be seamlessly integrated with several learning-based correspondence methods to handle complete-to-partial ambiguity. The proposed module is differentiable, fast, and parameter-free, as it only involves feature resampling and matching and runs in parallel on the GPU. The proposed module contributes to advancing the robustness and reliability of fully automatic end-to-end point cloud registration.
Limitations. The registration results depend on the condition that the source and target point clouds have similar densities. This may not hold if a very sparse target point cloud is manually collected via optical trackers. However, point cloud resampling methods may alleviate this limitation.
Experiments conducted in this paper are based on our constructed in silico and in vitro datasets. While our experiments have promisingly established a solid foundation for real-world applications, in future work, we will seek possibilities to construct an in vivo dataset to validate the performance of correspondence-based methods further.
VII Conclusion
We have constructed a benchmark using both in silico and in vitro datasets to evaluate state-of-the-art learning-based point cloud matching methods. Our analysis underscores the challenge of complete-to-partial ambiguity in scenarios with low visibility ratios, as typically occurs during liver surgery. In light of this view, we introduce a patches-to-partial strategy as a module that can be seamlessly integrated with learning-based point cloud correspondence methods.
We have demonstrated the effectiveness of this module in improving several methods, especially in low-visibility conditions. To summarize, the constructed benchmark and the proposed module lay a solid foundation for future studies of in vivo applications.